To treat a burn on your hand, you decide to place an ice cube on the burned skin. The mass of the ice cube is 19.7 g, and its initial temperature is -11.3 °C. The water resulting from the melted ice reaches the temperature of your skin, 29.0 °C. How much heat is absorbed by the ice cube and resulting water? Assume that all the water remains in your hand.
Answer #1

Given,

Mass of ice cube = 19.7 g

Initial Temp. = -11.3 degree C

Final Temp. = 29 degree C

The total heat absorbed in the above process can be broken down into 3 parts

1) Heat released when the ice is cooled from -11.3 to 0 degree C

2) Heat released when the ice melts at 0 degree C

3) Heat released when the temp. of water rises from 0 to 29 degree C

We know that,

Heat capacity for ice = 2.11 J / g K

Heat capacity for water= 4.18 J / g K

Heat of fusion for ice = 333.55 J / g

1) Heat released when the ice is cooled from -11.3 to 0 degree C

= 19.7 x 2.11 x (0 - (-11.3) = 469.71 J

2) Heat released when the ice melts at 0 degree C

= 19.7 x 333.55 = 6570.94 J

3) Heat released when the temp. of water rises from 0 to 29 degree C

= 19.7 x 4.18 x (29 - 0) = 2388.03 J

Total Heat absorbed by ice cube = 469.71 + 6570.94 + 2388.03 = 9428.68 J = 9.43 kJ